50 deg. C correspond to how many deg. F


1
Expert's answer
2020-07-20T06:23:34-0400

Temperature value can be expressed by three measurement units which include, degree Celsius("^0C" ), Kelvin(K) and degree Fahrenheit("^0F" ).A Fahrenheit temperature scale indicates that the melting point of water is 32 degrees and its boiling point 212 degrees. Conversions between Celsius and Fahrenheit are evaluated independently since they are not directly proportional.


The following formula is used to convert ("^0C" ) to ("^0F" )

"^0F=\ ^0C\ \ *\dfrac{9}{5}\ \ +32"


"= \ 50\ \ *\ \dfrac{9}{5}\ \ +32"

"=122^0\ F"